Use the following information for all 5 questions:  You manage a population of 75 finches on a farm. These finches can either have sharp beaks or wide beaks. This is a monogenic trait determined by a pair of alleles. The dominant allele, V, determines the sharp beak phenotype, while the recessive allele, v, determines the wide beak phenotype. The population contains 36 finches that are homozygous dominant, 15 that are heterozygotes, and 24 that are homozygous recessive.

What is the genotype of a “homozygous dominant” finch?

How many total alleles are in the gene pool, and how many of each allele are?

Calculate the allele frequencies for each.

Using the allele frequencies you calculated above, how many of each genotype would you have in a separate population of 100 finches, if the allele frequencies are the same? The Hardy-Weinberg Equation is p2 + 2pq + q2 = 1. Round to the nearest whole number
Answer #1

Genotype of homozygous dominant finch is VV

frequency of genotype = total number of individuals of that genotype / total number of individuals in population.

let p and q be the frequency of allele V and v

therefore frequency of genotype VV = 36 / 75 = 0.48 = p2

   frequency of genotype Vv = 15 / 75 = 0.20 = 2pq

   frequency of genotype vv = 24 / 75 = 0.32 = q2

therefore, frequency of allele V = p = 0.481/2 = 0.692

   frequency of allele v = q = 0.321/2 = 0.565

therefore total number of allele = frequency of that allele x total number of individuals in population

   total number of allele V = 0.692 x 75 = 52

   total number of allele v = 0.565 x 75 = 42

   total number of all alleles = 52 + 42 = 94

now , p2 + 2pq + q2 is approximately equal to 1 therefore frequencies of genotype and alleles will be same ( Hardy Weinberg equation )

therefore, in the separate population of 100 finches,

total number of genotype = frequency of that genotype x total number of individuals in population.

total number of genotype VV = 0.48 x 100 = 48

total number of genotype Vv = 0.20 x 100 = 20

total number of genotype vv = 0.32 x 100 = 32